OpenStudy (anonymous):

I need help understanding how to solve systems with matrices and how to do those constraints things. D;

OpenStudy (anonymous):

the constraints are \[ x \ge0\] \[ y \ge0\] \[y \le \frac{ 1 }{ 3 }x+3\] \[5\ge y+x\] and the Objective function is: C=6x-4y

OpenStudy (amistre64):

when the constraints are graphed, they form a critical region, an area that usually has corners where the lines meet. these corners represent critical points, one of them produces a max, another a min.

OpenStudy (anonymous):

@amistre64 what's a max/min ?

OpenStudy (amistre64):

a max/min is the maximum or minimum value that the objective function has over the area defined by the constraints

OpenStudy (amistre64):